find two consecutive odd integers such that 4 times the larger is 29 more than 3 times the smaller
Answer by ewatrrr(24785)


Hi,
I. Assigning the variable
Let x and (x+2) represent the smaller and larger of the two consecutive odd integers
II. Writing as we Read***
4(x+2) = 3x + 29
III. Solving for x
4x + 8 = 3x + 29
x = 21, smaller odd integer, 23, the larger of the two.
IV. CHECKING Our Answer***
4*23 = 3*21 + 29 = 92